OpenStudy (anonymous):

Find the slope-intercept form of the given line that passes through the points (-4, -1) and (3, -3)???? Show full work.

OpenStudy (anonymous):

Slope intercept form is the form \[y=mx+b\] where m is the slope and y is the y-intercept

OpenStudy (anonymous):

do you know how to find the slope between the 2 given points?

OpenStudy (anonymous):

No, unfortunately I don't.

OpenStudy (anonymous):

That's alright! So here's the formula... \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

So how would you plug in those points into this formula?

OpenStudy (anonymous):

-3-(-1)/3-(-4)??

OpenStudy (anonymous):

Exactly! :D

OpenStudy (anonymous):

That simplifies to -1/7

OpenStudy (anonymous):

So that's our m, now we gotta find the b

OpenStudy (anonymous):

And this is what we have so far... \[y=-\frac{ 1 }{ 7 }x+b\]

OpenStudy (anonymous):

Okay, so how do you find b?

OpenStudy (anonymous):

Hold on a sec, lemme check if I'm doing this right

OpenStudy (anonymous):

Ah! Okay - so all you gotta do now is just choose any one of the 2 points and plug the x and y values as well as the slope... so that the only variable left in our equation is b. Forget what I said about x=0, i was confusing it with something else

OpenStudy (anonymous):

It doesn't matter which point??

OpenStudy (anonymous):

Nope, both should give you the same answer

OpenStudy (sloppycanada):

Nope. But I recommend doing the easier one to solve. Usually it's positive.

OpenStudy (anonymous):

So, wait...-4=-1/7(-1)+b????

OpenStudy (anonymous):

Yeah! You can do that! :)

OpenStudy (anonymous):

But what is b??

OpenStudy (sloppycanada):

You solve for b. Like you're solving a regular equation.

OpenStudy (anonymous):

^^ Exactly

OpenStudy (sloppycanada):

So if I were you I'd factor out one seventh of -1.

OpenStudy (anonymous):

what?...is b a whole or mixed number...cause i got like -29/7 which seems weird. I'm not good at math so bear with me.

OpenStudy (anonymous):

\[-4=-\frac{ 1 }{ 7 }(-1) + b\] you can start by multiplying -1/7 and -1, you get \[-4=\frac{ 1 }{ 7 } +b\] then you take 1/7 to the other side and you get \[-4-\frac{ 1 }{ 7 }=b\] Can you take it on from here?

OpenStudy (anonymous):

i think so...

OpenStudy (anonymous):

you simply both sides and then isolate the b right?

OpenStudy (anonymous):

So it's \[b=-\frac{ 28 }{ 7 }-\frac{ 1 }{ 7 }=-\frac{ 29 }{ 7 }\] That make sense? I used common denominators.

OpenStudy (anonymous):

yeah that's what I got, -29/7. So I just leave it like that?

OpenStudy (anonymous):

I always prefer fractions instead of decimals, but it's up to you.

OpenStudy (anonymous):

Awesome, thanks so much for your help! :)

OpenStudy (anonymous):

Your welcome :)